Prepare a buffer solution with: 0.5 g NaC2H3O3 + 2 mL of 3 M acetic acid + 12 mL water Use 5 mL buffer + 0.2 mL 6 M NaOH What is the pH using the ICE table and Henderson Equation?
Answer #1

Ans. Step 1: Preparation of original buffer:

Moles of NaC2H3O2 = Mass / MW = 0.5 g / (82.034388 g/ mol) = 0.006095 mol

Note: Sodium acetate is CH­3COONa or NaC2H3O2 but NOT NaC2H3O3

# Total volume made upto = 2 mL (acetic acid) + 12 mL (water) = 14.0 mL

It’s assumed that addition of 0.5 g NaC2H3O2 does not affect the overall volume.

# Let denote-

            CH3COOH = AH                   ; CH3COO- = A-

# [AH] in final buffer solution:

Using C1V1 (original solution) = C2V2 (final buffer solution)

            Or, [AH], C2 = (3 M x 2.0 mL) / 14.0 mL = 0.4286 M

# [A-] in final buffer solution = Moles of NaC2H3O2 / Vol of buffer soln. in liters

                                                = 0.006095 mol / 0.014 L

                                                = 0.4354 M

# Step 2: Addition of NaOH to 5.0 mL original Buffer:

Total volume of solution = 5.0 mL + 0.2 mL = 5.2 mL

# [AH] = (0.4286 M x 5.0 mL) / 5.2 mL = 0.4121 M

[A-] = (0.4354 M x 5.0 mL) / 5.2 mL = 0.4187 M

[NaOH] = (6.0 M x 0.2 mL) / 5.2 mL = 0.2308 M

[H3O+] = 10-14 / [OH-] = 10-14 / 0.2308 = 4.33 x 10-14 M

# The two approaches of pH calculation are as follow-

Weak acid = AH = CH3COOH Conjugate base A CH3COC Given, [AH] in original Buffer = 0.4286 Origina!Vol of buffer = ;pKa-log Ka4

AH Weak acid Reaction: Given: Ka of AH1.80E-05 Create an ICE table as shown in fiqure- :A Conjugate base Given, [NaOH [OH- So